solve the equation by completing the square.
(x)square + 2x-2 = 0

Answer by Nate(3500)

x^2 + 2x - 2 = 0
x^2 + 2x = 2
(x + 1)^2 = 3
x + 1 = +- sqrt(3)
x = -1 +- sqrt(3)